Circular No. 6692

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 M1 (SOHO) C. St. Cyr, Naval Research Laboratory, reports, for the SOHO-LASCO

Consortium (see IAUC 6685), his discovery of another comet, fainter than m_{1} = 5, in SOHO-LASCO C3 coronagraph images. The object was somewhat elongated along the sun-comet line, and, as with many of the other Kreutz members found by SOHO, he considers this elongation to be a faint tail. Measurements by St. Cyr and S. Stezelberger have been reduced by G. V. Williams and are given on MPEC 1997-N04 in detail, the approximate discovery position being as follows:

1997 UT  α_{2000}  δ_{2000}
June 29.801  6^{h}24.^{m}0  +19^{o}24′

Orbital computations by the undersigned, also given in detail on MPEC 1997-N04, confirm the comet’s membership in the Kreutz group and indicate perihelion passage on June 30.95.

COMET 1997 N1 Vello Tabur reports his visual discovery of a comet, for which the

following observations are available:

1997 UT  α_{2000}  δ_{2000}  m_{1} Observer
July 2.7986 4^{h}48.^{m}3  −26^{o}58′  10 Tabur
2.8257  4 48.5  −26 58  _{”}
3.79386 4 55 30.^{s}93 −26 40 16′′.4  Garradd
3.79570 4 55 31.73 −26 40 14.3  _{”}
3.79820 4 55 32.81 −26 40 11.6  _{”}
3.80794 4 55 37.09 −26 40 00.4  _{”}
3.81361 4 55 39.57 −26 39 53.8  _{”}
3.82903 4 55 46.41 −26 39 35.3  Child
3.83319 4 55 48.17 −26 39 30.2  _{”}
3.83633 4 55 49.47 −26 39 27.8  _{”}
3.84062 4 55 51.28 −26 39 23.3  _{”}
3.84645 4 55 53.93 −26 39 16.4  _{”}

V. Tabur (Wanniassa, A.C.T.). 0.20-m f/4.7 reflector. Rather uncon-

densed circular coma ∼ 1′.5 across, no tail. G. J. Garradd (Loomberah). 0.25-m f/4.1 reflector + CCD. A 60-s 2 × 2

binned exposure shows a 2′.5 tail in p.a. 210^{o}. J. B. Child (Macquarie). 0.32-m f/4.8 reflector + CCD. 3′ tail.

1997 July 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden